OpenStudy (anonymous):

Can someone help me prove this ( I would type the ques ASAP) @satellite73

OpenStudy (anonymous):

If a and b be two real roots of the equation x^3 +px^2 + qx +r = 0 satisfying the relation ab+1=0 , then prove that r^2 +pr + q +1 = 0

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

So, considering c as the third root \[\huge a + b + c = -p \] \[\huge ab +ac + bc = q \] \[\huge abc = -r \]

OpenStudy (anonymous):

i got no idea, but doesn't the last equation tell you \(c=r\) ?

OpenStudy (anonymous):

yes it does i am working , i will notify you when i get the answer

OpenStudy (anonymous):

maybe that is all you need

OpenStudy (anonymous):

factor the \(r\) out of \[r^3+pr^2+qr+r=0\]

OpenStudy (anonymous):

you get \[r^2+pr+q+1=0\]

OpenStudy (anonymous):

Wait i think i got it

OpenStudy (anonymous):

you do

OpenStudy (anonymous):

once you have \(r\) is a root, then it all works out nicely

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

We get c = r and c(b+a) =q+1 So, we have to prove r^2 +pr + q +1 =0 c^2 -(a+b+c)c +bc+ac c^2-ac -bc -c^2 +bc +Ac 

OpenStudy (anonymous):

i think you are doing too much work

OpenStudy (anonymous):

you already said the third real root is \(r\) right? because \[abc=-r\] and \(ab=-1\) making \(c=r\) so \(r\) is a root

OpenStudy (anonymous):

that means \[r^3+pr^2+qr+r=0\]

OpenStudy (anonymous):

which also tells you \[r(r^2+pr+q+1)=0\] and since \(r\neq 0\) that makes \[r^2+pr+q+1=0\]

OpenStudy (anonymous):

nice!

OpenStudy (anonymous):

Actually i have got little confidence doing these proving sums

OpenStudy (anonymous):

you did all the work it came down to figuring out that the third root was \(r\)

OpenStudy (anonymous):

yes that's why i am so excited